OpenStudy (anonymous):

write the equation for the line containing (2,4) and perpendicular to y=2x-3.

OpenStudy (anonymous):

m = -1/2 y - 4 = -1/2 ( x - 2) 8 - 2y = x - 2 x + 2y = 10

OpenStudy (anonymous):

that wrong

OpenStudy (anonymous):

there 2 option y=1/2 + 3 or y=1/2+5

OpenStudy (anonymous):

its y=-1/2 +5 he was right, but u have to isolate the y and perpendicular lines are the opposite (negative/positive) of the reciprical (idk how u spell it)